OpenStudy (anonymous):

All of the following points lie on the graph of y = 2x(the x is an exponent)except _____. (0, 0) (1, 2) (2, 4) (3, 8)

OpenStudy (anonymous):

Substitite x and y coordinates of every point for respectively x and y in the equation and see in which case it isn't equal.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

I'm new to this stuff so... no.

OpenStudy (anonymous):

Oh, ok. An example: let's take a point (1,2). Its x-coordinate is 1 and its y-coordinate is 2. If it lies on our line, then these numbers must solve (so give 0=0 in the end) its equation. Let's see: 2 = 2^1 2 = 2 0 = 0 Ok, so this point indeed lies on the line. Do the same stuff with other points, ok?

OpenStudy (anonymous):

Alright.. Thank you~

OpenStudy (anonymous):

To make myself sure - what answer did you get? :)

OpenStudy (anonymous):

I'm trying to do the problem right now.. >^<

OpenStudy (anonymous):

Oh, sorry for disturbing then.

OpenStudy (anonymous):

Is it (3,8) ? >^<

OpenStudy (anonymous):

Unfortunately no. Because then x=3 and y=8. Is it true that 8 = 2^3? It is, so it must lie on the line. Take a look at the point (0,0), though. What is 2^0?

OpenStudy (anonymous):

*sighes* I think I need to hit the books. I don't understand this at all. I am absolutely terrible at math.

OpenStudy (anonymous):

It must be a minor thing that's blocking you. Okay, so I'll show you how to do this at least. For point (0,0), we have x=0 and y=0. Let's substitute this into the equation of our line: y = 2^x 0 = 2^0 But 2^0 = 1. (Perhaps you didn't know that?) We get 0=1, but it is not true! Therefore this point doesn't lie on the line, because its coordinates doesn't satisfy the equation.